OpenStudy (newwar):

What is 4.6 × 108 cm/min in meters per hour? A. 2.76 × 108 m/h B. 2.76 × 1010 m/h C. 7.67 × 106 m/h D. 7.67 × 108 m/h

OpenStudy (newwar):

@FitzIncorporated

OpenStudy (anonymous):

\(4.6 \times 10^8 \) \(cm/\min\)  therefore, \(\huge (\frac{ 4.6 }{ 100 }\times 60)\) \(\huge x\) \(\huge 10^8\) \(\huge m/hr\)